OpenStudy (anonymous):

The diameter of the circle above is 94 cm. What is the circumference of the circle? (Use = 3.14.) A. 295.16 cm B. 13,872.52 cm C. 6,936.26 cm D. 147.58 cm

OpenStudy (shamil98):

Cicumference = 2piR 2R = D 94/2 = r Can you do it from here?

OpenStudy (anonymous):

no i am so stuck and i can try thought thanks

OpenStudy (shamil98):

C=PiD as well

OpenStudy (shamil98):

So C = 3.14(94)

OpenStudy (shamil98):

Just multiply the values to get your circumeference

OpenStudy (shamil98):

\[\large C= \Pi D\] \[\large C = 3.14(94)\]

OpenStudy (anonymous):

so C is the answer?

OpenStudy (shamil98):

No. C stands for circumeference..

OpenStudy (shamil98):

What is 94 times 3.14? use your calculator

OpenStudy (anonymous):

295.16

OpenStudy (anonymous):

so that is the answer

OpenStudy (shamil98):

Yes it is good job :)

OpenStudy (anonymous):

yay thanks